Please help me solve this math problem:
Tom is 3 times as old as Todd. In 4 years Tom's age will be 6 years more than 2 times as old as Todd. How old is Todd?

I've tryed using a Chart to slove this math problem. But it wasn't working out. Can you help me find an easyer way to slove it?

Found 2 solutions by scott8148, Earlsdon:
Answer by scott8148(5895)

let x=Todd, y=Tom

now ... y=3x

4 years from now ... y+4=6+2(x+4) ... y+4=6+2x+8 ... y=2x+10

substituting gives 3x=2x+10 ... x=10

Answer by Earlsdon(6103)

Try this:
Let Tm = Tom's present age and Td = Todd's present age.
From the problem description you can write:
1) Tm = 3*Td "Tom is 3 times as old as Todd"
2) Tm + 4 = 2(Td+4)+6 "In 4 years Tom's age will be 6 years more than 2 times as old as Todd".
Substitute Tm = 3*Td from equation 1) into equation 2) and solve for Td.
2a) 3*Td +4 = 2(Td+4)+6 Simplify.
3Td+4 = 2Td+8+6
3Td+4 = 2Td+14 Subtract 2Td from both sides.
Td+4 = 14 Subtract 4 from both sides.
Td = 10 This is Todd's present age.